OpenStudy (anonymous):

plants of a particular species have a 25% chance of being red-flowering, independently of other plants. Find the normal approximation to the chance that among 10,000 plants of this species, more than 2400 are red-flowering.

OpenStudy (anonymous):

@sara98

OpenStudy (anonymous):

@benjamin2123

OpenStudy (anonymous):

@pandu

OpenStudy (kropot72):

\[mean=np=10,000\times 0.25=2500\] \[\sigma=\sqrt{10,000\times 0.25\times 0.75}=43.3\] The z-score for 2400 is given by \[z=\frac{X-\mu}{\sigma}=\frac{2400-2500}{43.3}=-2.31\] Now you need to find the required probability by referring to a standard normal distribution table.

OpenStudy (anonymous):

THAT WOULD BE .990

OpenStudy (anonymous):

WHICH IS WRONG BY THE WAY

OpenStudy (kropot72):

I have recalculated the z-score using the continuity correction of 0.5. The z-score then becomes -2.3 leading to the result that there is a 98.93% chance that more than 2400 will be red-flowering.

OpenStudy (anonymous):

WOW .01 MADE SUCH A BIG DIFFERENCE

OpenStudy (anonymous):

SO 0.9893 IS THE RIGHT ANSWER

OpenStudy (kropot72):

I wish that I used the continuity correction the first time. Sorry about that!